# How many simulations do we need for simulation-based inference in cosmology?

### v1

@@ V.1 The Big Sobol Sequence (BSQ) Simulations
Figure 4 shows the relationship between loss and the number of simulations, demonstrating that with sufficient training data, our neural network achieves optimal summary statistics as the loss curve approaches the theoretical information bound. The BSQ simulations exhibit a slightly higher test loss compared to the LH simulations. However, the test loss follows a power law scaling of $N^{-0.36}$, remarkably consistent with our earlier prediction from the Latin Hypercube simulations, validating our methodology within 10% accuracy.
@@
The parameter uncertainties inferred from models trained on different numbers of simulations appear in Figure 5. These results demonstrate that the information content saturates nearly 4,000 simulations, with minimal improvement beyond this point. This finding validates our earlier prediction based on power law scaling analysis of the Latin Hypercube simulations, confirming that approximately 4,000 simulations are required to extract the full information content from $P(k)$.
@@
Figure 4: Test loss (IV.1) as a function of number of BSQ simulations. The loss follows a simple power law scaling law across 2 orders of magnitude with a power law similar to the LH simulations. The combined test loss for all parameters does not reach the optimal Fisher information computed at the fiducial point but covers the full prior range in parameter space. The loss contributions of BSQ simulations are slightly higher than the LH simulations, but Figure 5 shows that the NN achieves asymptotic optimality at the fiducial point at 4000 simulations, as predicted.
@@
VI Discussion and Conclusion
This paper addresses a fundamental question in simulation-based inference for cosmology: How many simulations are required for neural networks to achieve optimal parameter inference? Our investigation of this question yields several important insights.
First, we demonstrate that current simulation suites, while impressive in scope, may be insufficient for optimal neural network training. Specifically, we show that the 2,000 Latin Hypercube simulations in the Quijote suite, while valuable, do not provide enough training data for neural networks to reach the Cramér-Rao information bound, even for the relatively simple case of the matter power spectrum $P(k)$.
@@ VI Discussion and Conclusion
This prediction motivated the creation of our Big Sobol Sequence (BSQ), a suite of 32,768 N-body simulations that allows us to verify these scaling relationships and provide a comprehensive resource for future studies. Analysis of networks trained on the BSQ simulations confirms our earlier predictions. The test loss follows a power law decay similar to that observed in the Latin Hypercube simulations, with information content indeed beginning to saturate around 4,000 simulations. This consistency across different simulation designs suggests we have identified a fundamental aspect of neural network learning in cosmological parameter inference. Our architectural studies reveal an important relationship between network depth and training set size. While deeper networks can potentially extract more information, we find that increasing the number of training simulations often provides greater improvement than adding network layers.
@@
Looking forward, several questions merit further investigation. The power law scaling we observe may have deeper theoretical foundations that could inform future network design. Additionally, extending this analysis to other summary statistics and their combinations could provide valuable insights for survey analysis. The BSQ simulation suite, which we make publicly available, provides a robust foundation for such investigations. Current studies report field-level inference on weak lensing data outperforms when the NN training is conditioned on some complimentary summary i.e. power spectrum $P(k)$; this method provides optimal information when there are not enough simulations for training[49]. Although similar analysis on dark matter halo fields shows no improvement for 2000 Latin hypercube simulations. In our upcoming paper, we show a patch-based estimator to get optimal information from any survey volume without generating training simulations of present-day survey volume or requiring huge computational resources to train such a large high-resolution volume.
The methods and insights presented here extend beyond power spectrum analysis. As we prepare for next-generation surveys, understanding the requirements for optimal neural network performance becomes increasingly crucial. Our work provides both practical guidance for simulation campaigns and a framework for guiding neural information extraction in cosmological parameter inference.
VII Acknowledgements
@@
VIII Appendix
VIII.1 Effect of varying the network architecture
@@
Figure 6: A generic MLP architecture of varying number of hidden layers of dimension 128. This is used to analyze the impact of the depth of the NN on the inference.
@@
Given the computational cost of hyperparameter optimization for deep architectures, we conducted an extensive hyperparameter search using a single-layer perceptron (equivalent to linear regression) on $P(k)$. This approach aligns with recent studies of large neural networks[47], which suggest hyperparameter scalability across network depths. We utilize the Weights and Biases[50] tool to optimize key hyperparameters for our neural network. Our search explores batch sizes in the range $[16,64]$, activation functions (tanh, ReLU, sigmoid, and a custom activation function (see appendix VIII.2)), optimizers (SGD, Adam, and BFGS), and learning rate schedulers (CyclicLR and StepLR). Additionally, we examine learning rates within $[0,0.01]$ and tune momentum and gamma values in the range $[0.1,1]$. This study aims to identify the optimal combination of hyperparameters to enhance model performance effectively. The optimal hyperparameter values appear in Table 1.
@@
VIII.2 Activation Function

### v2

@@ 5.1 The Big Sobol Sequence (BSQ) simulations
Figure 4 shows the relationship between loss and the number of simulations, demonstrating that with sufficient training data, our neural network achieves optimal summary statistics as the loss curve approaches the theoretical information bound. Networks trained on the BSQ simulations exhibit a slightly higher test loss compared to the LH simulations for the same size of training set. The slightly improved accuracy of the BSQ simulations is unlikely to produce a noticeable effect on the power spectra in the $k$ range used for this study. A more likely reason is the different sampling of the parameter space of the Sobol sequence and the Latin Hypercube designs. This difference in sampling could lead to variations in the effective training set size and the resulting model performance. However, the test loss of BSQ simulations does also follow a power law scaling of $N^{-0.36}$, with a power law exponent that is within 10% of the exponent we found earlier from the Latin Hypercube simulations.
@@
The parameter uncertainties inferred from models trained on different numbers of simulations appear in Figure 5. These results demonstrate that the information content saturates for nearly 4,000 simulations, with minimal improvement beyond this point. This is consistent with our prediction based on the power law scaling analysis of the Latin Hypercube simulations.
The Fisher matrix results seem to be better than expected based on the higher test loss in the BSQ suite. Based on the offset of the scaling laws we may have expected that more, perhaps 10,000 simulations would have been necessary to reach the asymptotic regime. Consider that the Fisher information is computed at the fiducial model in the center of the parameter space while the test loss is an average over the full prior volume. The Sobol sequence is strongly self-avoiding and so there will be models in the test set that are near the boundaries of the prior space, where the feature extraction may become less efficient. In summary, we find that the difference in parameter sampling do affect the constant in the test loss scaling; the two sampling designs seem to provide similar Fisher information for similar training set size at the fiducial parameter set in the center of the prior volume.
@@
Figure 4: Test loss (5) as a function of number of BSQ simulations. The loss follows a simple power law scaling law across 2 orders of magnitude with a power law slope similar to the LH simulations (shown for comparison). The combined test loss for all parameters does not reach the optimal Fisher information computed at the fiducial point. The test loss is computed over the full prior range in parameter space rather than at a single point in the center of the training data. The loss of BSQ simulations are slightly higher than the LH simulations, but Figure 5 shows that the NN achieves asymptotic optimality at the fiducial point at 4000 simulations.
@@
5.2 Robustness of scaling laws: Wavelet statistics
So far, we collected evidence for scaling laws that for feature extraction from the power spectrum using a quadratic loss function. In this subsection, we will go beyond this and look for scaling laws in a simulation based inference task based on Wavelet Scattering Transform (WST) coefficients (Valogiannis & Dvorkin 2022; Eickenberg et al. 2022; Blancard et al. 2024; Gatti et al. 2024).
We have computed our wavelet statistics using the Kymatio package (Andreux et al. 2022) from Quijote dark matter simulations of $128^{3}$ down-sampled resolution. To reduce the dynamic range we take the natural log of the WST coefficients as the summary vector. We find that some WST coefficients are highly correlated in the fiducial simulations; this causes difficulties downstream, since such correlations render the covariance matrix nearly singular. This in turn leads to numerical instabilities in the Fisher matrix computation. To address this, we find highly correlated pairs (those with Pearson’s $|r|>0.99$ (Kendall & Stuart 1977)) and then eliminate redundant WST coefficients from the parameter vector.
We use Learning the Universe Implicit Likelihood Inference (LtU-ILI) (Ho et al. 2024) with the lampe backend to train an ensemble of neural posterior estimators (NPEs) consisting of Masked Autoregressive Flow (MAF) (Papamakarios et al. 2017), Gaussianization Flow (gf) (Meng et al. 2020), Nonlinear Independent Components Estimation (NICE) (Dinh et al. 2014), Neural Spline Flow (NSF) (Durkan et al. 2019), and Mixture Density Network (MDN) (Bishop 1994) on the wavelets and estimate the log-probability of the fiducial simulations. We repeat this process for each number of training simulations. The results suggest that the log-probability again follows a scaling law as the number of training simulations changes, as shown in Figure 6.
Figure 6: Log-probability from an ensemble of $N$ trained on wavelets for different number of Quijote LH training simulations. The log-probability simply follows a power law similar to the $P(k)$ analysis.
We draw 10000 samples from the learned neural posteriors for a simulation at the fiducial cosmology and use those samples to demonstrate the posterior distribution in Figure 7 based on training with 100, 500, 1000 and 1500 simulations. The corner plot shows that the constraints on $\Omega_{m}$ and $\sigma_{8}$ decrease significantly with the increasing number of simulations, while other parameters show mild improvement with the number of training simulations.
Figure 7: Scaling of information content of the neural posterior estimation based on wavelet summaries. The power law scaling of the log probability loss with the increasing number of simulations is reflected in improved constraints on $\Omega_{m}$ and $\sigma_{8}$. Other parameters show mild improvement.
6 Discussion and conclusion
This paper addresses a fundamental question in simulation-based inference for cosmology: How many simulations are required for neural networks to achieve optimal feature extraction for parameter inference? Our investigation of this question yields several important insights.
First, we demonstrate that current simulation suites, while impressive in scope, may be insufficient for optimal training of neural feature extractors. Specifically, we show that the 2,000 Latin Hypercube simulations in the Quijote suite, while valuable, do not provide enough training data for neural networks to reach the Cramér-Rao information bound, even for the relatively simple case of the matter power spectrum $P(k)$.
@@ 6 Discussion and conclusion
This prediction motivated the creation of our Big Sobol Sequence (BSQ) [^3] https://quijote-simulations.readthedocs.io/en/latest/bsq.html, a new and publicly available suite of 32,768 N-body simulations that allows us to verify these scaling relationships and provide a comprehensive resource for future studies.
Analysis of networks trained on the BSQ simulations confirms our earlier predictions. The test loss follows a power law decay similar to that observed in the Latin Hypercube simulations, with information content indeed beginning to saturate around 4,000 simulations. While the sample efficiency of the two simulation designs differs, producing different constant factors for the scaling laws, each simulation suite exhibited predictable power law scaling of the test loss with the size of the training set.
Extending the analysis beyond the power spectrum and squared error loss, we ran a similar scaling study for neural posterior estimation of cosmological parameters from wavelet scattering transform coefficients. We find that the test Kullback-Leibler loss for the neural density estimation task also scales as a power law of training set size.
Our architectural studies reveal an important relationship between network depth and training set size. While deeper networks can potentially extract more information, we find that increasing the number of training simulations often provides greater improvement than adding network layers.
@@
Looking forward, several questions merit further investigation. The power law scaling we observe may have deeper theoretical foundations that could inform future network design. The difference we observe in sampling efficiency between the Sobol sequence and the Latin Hypercube sampling designs suggests that surveying other strategies to sample parameters for training sets may lead to more simulation efficient information extraction. Additionally, extending this analysis to other summary statistics and their combinations could provide valuable insights for survey analysis. The BSQ simulation suite, which we make publicly available, provides a robust foundation for such investigations.
Our study also motivates research into more simulation efficient ways to train neural networks for feature extraction. A recent study reports that in the case of limited training data, machine learning inference on weak lensing data outperforms when the NN training is conditioned on some complementary summary i.e. power spectrum $P(k)$ (Makinen et al. 2024). The steeper scaling law we found in our case study of neural posterior estimation compared to estimating the feature extraction using regression of the posterior mean also suggests that some inference approaches may be more sample efficient than others. In the interest of reducing the simulation cost to train feature extractors, we plan to investigate hierarchical approaches that avoid training networks that ingest the full survey volume at full resolution all at once.
As we prepare for next-generation surveys with larger volumes and higher galaxy number density, understanding the requirements for neural network performance becomes increasingly crucial. Our work provides both practical guidance for simulation campaigns and a framework for guiding neural information extraction in cosmological parameter inference.
Acknowledgements.
@@
Appendix A Effect of varying the neural network architecture
@@
Figure 8: A generic MLP architecture of varying numbers of hidden layers of dimension 128. This is used to analyze the impact of the depth of the NN on the inference.
@@
Given the computational cost of hyperparameter optimization for deep architectures, we conducted an extensive hyperparameter search using a single-layer perceptron (equivalent to linear regression) on $P(k)$. This approach aligns with recent studies of large neural networks (Kaplan et al. 2020), which suggest hyperparameter scalability across network depths. We utilize the Weights and Biases (Biewald 2020) tool to optimize key hyperparameters for our neural network. Our search explores batch sizes in the range $[16,64]$, activation functions (tanh, ReLU, sigmoid, and a custom activation function (see appendix B)), optimizers (SGD, Adam, and BFGS), and learning rate schedulers (CyclicLR and StepLR). Additionally, we examine learning rates within $[0,0.01]$ and tune momentum and gamma values in the range $[0.1,1]$. This study aims to identify the optimal combination of hyperparameters to enhance model performance effectively. The optimal hyperparameter values appear in Table 1.
@@
Appendix B Activation functions